40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jan 28 18:03:53 2021, shruti, DailyProgress, PSOMA, cavity locking Path.pdf
    Reply  Mon Feb 1 18:25:47 2021, shruti, DailyProgress, PSOMA, cavity locking 
       Reply  Wed Feb 3 10:56:54 2021, shruti, Update, PSOMA, cavity locking 
          Reply  Thu Feb 4 16:19:35 2021, shruti, Update, PSOMA, cavity alignment 5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg
             Reply  Thu Feb 4 22:27:43 2021, rana, Update, PSOMA, cavity alignment IMC_HOManalysis.ipynb
                Reply  Tue Feb 9 11:14:48 2021, shruti, Update, PSOMA, cavity alignment 
                   Reply  Tue Feb 16 11:30:14 2021, shruti, Update, PSOMA, Beat recovered, fiber organization, alignment F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg069D0D53-9111-4B33-96AD-05B907C4BBCB.jpegBeat_low_freq_oscillations.mp4
Message ID: 2630     Entry time: Thu Feb 4 16:19:35 2021     In reply to: 2626     Reply to this: 2631
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: cavity alignment 

I turned on both the lasers and TECs, and cleared out all extra optics/components from around the experiment.

When I re-checked the cavity alignment I saw a large beam off the side of the main beam when the main beam was not blocked, both on the third side (Mc -> Ma). I tweaked around with the alignment quite a bit and even though the beams seemed to overlap along one side (Ma -> Mb). I could not really check the third side properly.

The transmission PD shows that it is receiving power from more than one round trip as checked by blocking the outgoing light from optic Mc, but I'm still not sure if there is interference.

I also measured the beam profiles at different points on the path (will add the data later). but in summary,  the beam seems more or less collimated everywhere with a beam diameter of 2.4 mm.

 

[Rana, Shruti]

The fiber launch PAF2A-18C seems to have an output waist of 3.49 mm. We are ordering fiber launches with smaller waists, but in the meantime it was previously adjusted to have minimum divergence. Today Rana adjusted the three tip-tilt/ focus knobs to get the smallest waist before the entering the cavity, but it turned out result in the huge divergence within the cavity, then it was re-adjusted to a larger beam with lesser divergence, close to what it was initially.

We also experimented with a lens, but then decided against installing one after seeing that it did not help very much.

The transmission PD reading may be noisy/ somewhat incorrect because of scattered visible light.

QUESTION: Do we have filters for 1550 nm that blocks shorter wavelengths well?

Attachment 1: 5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg  3.107 MB  | Hide | Hide all
5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg
ELOG V3.1.3-